Solve the quadratic equation 6x^z - 11x - 7 = 0
Answer by scott8148(6628)

(3z - 7) (2z + 1) = 0

3z - 7 = 0 ___ 3z = 7 ___ z = 7/3

2z + 1 = 0 ___ 2z = -1 ___ z = -1/2